OpenStudy (mwhit2101):

SIXTH GRADE MATH HELP MEDAL+FAN! The cost of 4 kilograms of cheese is 50 dollars. What is an equation to show this, but in function notation?

OpenStudy (mathstudent55):

Let's call the cost y and weight x.

OpenStudy (mwhit2101):

Sure, ok.

OpenStudy (anonymous):

$12.5 per kilogram or \[1.25\times 10^{1}\]

OpenStudy (mathstudent55):

$50 corresponds to 4 kg That means there must be a number that when you multiply by 4 you get 50. Let that number be k. k is the price of cheese per kg. 1 kg: y = k * 1 2 kg: y = k * 2

OpenStudy (anonymous):

per kilogram

OpenStudy (mathstudent55):

3 kg: y = k * 3 4 kg: y = k * 4 For x kilograms: y = k * x

OpenStudy (mathstudent55):

Now let's look at 4 kg: y = k * 4, but the cost is 50, so we have 50 = k * 4 We can solve for k, the cost per kilogram, and get k = 12.5

OpenStudy (mathstudent55):

Now that we know what the cost per kg is, we can rewrite our function using the cost: y = 12.5x Since we want function notation, we finally write: f(x) = 12.5x

OpenStudy (mwhit2101):

None of my answer selections have anything to do with the number 12.5. Here are my selections: Cost(50)=4 Cost(4)=50

OpenStudy (mwhit2101):

@mathstudent55

OpenStudy (mwhit2101):

I think it's going to be the cost(4)=50. The cost depends on the number of cheeses...4. The cost is 50. I'm going to go with that one. Thanks anyway lol

OpenStudy (mathstudent55):

What are the other choices?

OpenStudy (mathstudent55):

Oh, I see. With only those two choices, you are correct. Cost is a function of weight. cost(4) = 50 makes sense.